OpenStudy (anonymous):

16 out of 25 . IS WHAT LETTER GRADE ?

OpenStudy (anonymous):

It is a 64. Usually an F depending where you go to school at.

OpenStudy (anonymous):

64 percent

OpenStudy (anonymous):

she's correct, its an F :/

hero (hero):

OMG, did someone get an F

OpenStudy (anonymous):

Actually thats a low D

OpenStudy (anonymous):

to find any letter grade, take out your calculator and divide. so in your case, you divide 16 by 25

OpenStudy (anonymous):

naw its a c+

OpenStudy (anonymous):

no, a c is at least a 70 %

OpenStudy (anonymous):

Depending upon where you live and where you attend school.

OpenStudy (anonymous):

no a c -= my homeschooling thing is different thats why

OpenStudy (anonymous):

A= 100-90 B= 89-80 C= 79-70 D= 69-60 F= 59 and below keep in mind some places do it slightly different, but as a general rule, thats it.

OpenStudy (anonymous):

I understand Lissaa_bih. I was homeschooled last year. Hang in there :/

OpenStudy (anonymous):

yeaah , but this homeschooling thing im doing is a whooole different thing

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

like everything is diffreent the grading levels the work all of it lol

OpenStudy (anonymous):

like everything is diffreent the grading levels the work all of it lol